OpenStudy (anonymous):

A pyramid has a volume of 18 in3. If all measures of the pyramid are multiplied by 6, what will the volume be?

OpenStudy (amistre64):

how do we determine the volume of a pyrimad?

OpenStudy (anonymous):

LxWxH then divide by 3.

OpenStudy (amistre64):

yep so each part is times 6 bigger l h w/3 = 18 l h w = 3(18) if we times the other side by the appropriate value/s ... what do we do to the other side?

OpenStudy (amistre64):

moving the 3 wasnt needed, but i seem to have felt like doing it anyways .... well move it back later

OpenStudy (anonymous):

I dont understand what you're asking.

OpenStudy (amistre64):

apply the given information, all measures of the pyramid are multiplied by 6 resulting in 6L, 6W and 6H whats the new volume?

OpenStudy (amistre64):

or think about it in this way: since we know that LWH/3 = 18 6L 6W 6H/3 is our new volume 6^3 LWH/3 = new volume, but LWH/3 is 18 6^3 18 = new volume